# Cryptographic Characterization of Quantum Advantage

### v1

@@ 2.3 One-Way Puzzles
It is known that $c$-correct and $s$-secure OWPuzzs with $c(\lambda)-s(\lambda)\geq 1/{\mathrm{poly}}(\lambda)$ can be amplified to $(1-{\mathsf{negl}}(\lambda))$-correct and ${\mathsf{negl}}(\lambda)$-secure OWPuzzs [CGG24b]. On the other hand, we do not know how to amplify the gap of classically-secure OWPuzzs.
@@ 3.4 QASs Imply Non-Interactive IV-PoQ
$\mathsf{SD}(\mathcal{A}(1^{\lambda}),\mathcal{B}(1^{\lambda}))\leq\frac{1}{N^{1/4}(\lambda)}\leq\frac{1}{q(\lambda)}$ (50)
@@
holds for all sufficiently large $\lambda\in\Lambda$, where $\delta\coloneqq\frac{1}{\lambda^{\log\lambda}}+\frac{1}{\sqrt{N(\lambda)}}$, which is less than $1/2$ for sufficiently large $\lambda$. In Equation 81, we have used the following lemma.
@@
$\mathsf{SD}(\mathcal{A}(1^{\lambda}),\mathcal{B}(1^{\lambda}))\leq\frac{1}{\lambda^{\log\lambda}}+\frac{1}{\sqrt{N}}+\sqrt{\frac{2\log^{2}\lambda}{N}}\leq\frac{N^{\frac{1}{4}}}{\sqrt{N}}=\frac{1}{N^{\frac{1}{4}}}$ (102)
@@ 5 Equivalence of IV-PoQ and Classically-Secure OWPuzzs
For this specific $\{f_{\lambda}\}_{\lambda\in\mathbb{N}}$, there exists a PPT algorithm $\mathcal{R}_{p}$ such that Equations 127 and 128 hold for infinitely many $\lambda\in\mathbb{N}$. We write $\Lambda\subseteq\mathbb{N}$ to mean the set of such $\lambda$. Rewriting Equation 128, we have
@@
As we have stated above, for this specific $\{f_{p,\lambda}\}_{\lambda\in\mathbb{N}}$, there exists a PPT algorithm $\mathcal{R}_{p}^{f_{p,\lambda}}$ such that Equations 225 and 226 hold for infinitely many $\lambda\in\mathbb{N}$. From $\mathcal{S}_{p}$ and $\mathcal{R}_{p}^{f_{p,\lambda}}$, we construct a PPT algorithm $\mathcal{A}$ such that
@@
holds for infinitely many $\lambda\in\mathbb{N}$. Here, we have used Equation 225 to obtain Equations 234 and 240, and used Equation 226 to obtain Equation 238. Therefore, $\mathcal{A}$ breaks the soundness of classically-secure OWPuzz $(\mathsf{Samp},\mathsf{Ver})$.
@@ 6.2 Zero-Knowledge IV-PoQ
We also do not have solution for the latter approach either, but we have the following observation. We observe that we can construct statistical zero-knowledge IV-PoQ (or even efficiently-verifiable PoQ) if we additionally assume the existence of an $\mathbf{NP}$ search problem that is easy for QPT algorithms but hard for PPT algorithms (or equivalently publicly-verifiable one-round PoQ). To see this, we can consider a protocol where the honest quantum prover solves the $\mathbf{NP}$ search problem and then proves the knowledge of the solution by using statistical zero-knowledge arguments of knowledge for $\mathbf{NP}$, which exists if OWFs exist [HNO+09]. Examples of classically-hard and quantumly-easy $\mathbf{NP}$ search problems are the factoring and discrete-logarithm problems (assuming classical hardness of them) [Sho94]. Another example based on a random oracle was recently found in [YZ24]. Thus, based on the random oracle heuristic [BR93], we have a candidate construction of statistical zero-knowledge PoQ from hash functions. [^33] This is not a construction in the quantum random oracle model since we use the hash function in a non-black-box manner. Instead, we rely on the assumption that the problem considered in [YZ24] is classically hard when the random oracle is instantiated with a concrete hash function. Though this is far from a construction solely based on OWFs, this can be seen as an evidence that “structured” assumptions are not necessary for statistical zero-knowledge PoQ, let alone for statistical zero-knowledge IV-PoQ.343434“Structure” is a commonly used informal term that refers to problems behind constructions of existing public key encryption such as the hardness of factoring, discrete-logarithm, learning with errors, etc. On the other hand, hash functions are often regarded as “unstructured”. See [Bar17] for more context.
@@ Appendix A On Uniformity of Adversaries
The time to compute $f^{\prime}(z)$ is $O(|z|^{2})$ because finding the maximum $i$ spends $O(|z|^{2})$ time363636 A single-tape Turing machine spends $O(|z|^{2})$ time to find the maximum $i$. and computing $f(z_{1,...,i})$ spends $O(|z_{1,...,i}|^{c})\leq O(|z|)$ time.

### v2

@@ 2.3 One-Way Puzzles
It is known that $c$-correct and $s$-secure OWPuzzs with $c(\lambda)-s(\lambda)\geq 1/{\mathrm{poly}}(\lambda)$ can be amplified to $(1-{\mathsf{negl}}(\lambda))$-correct and ${\mathsf{negl}}(\lambda)$-secure OWPuzzs [CGG24]. On the other hand, we do not know how to amplify the gap of classically-secure OWPuzzs.
@@ 3.4 QASs Imply Non-Interactive IV-PoQ
holds for all sufficiently large $\lambda\in\Lambda$, where $\delta\coloneqq\frac{1}{\lambda^{\log\lambda}}+\frac{1}{q(\lambda)^{2}}$, which is less than $1/2$ for sufficiently large $\lambda$. In Equation 81, we have used the following lemma.
@@
$\mathsf{SD}(\mathcal{A}(1^{\lambda}),\mathcal{B}(1^{\lambda}))\leq\frac{1}{\lambda^{\log\lambda}}+\frac{1}{q(\lambda)^{2}}+\sqrt{\frac{2\log^{2}\lambda}{N}}\leq\frac{q(\lambda)}{q(\lambda)^{2}}=\frac{1}{q(\lambda)}$ (102)
@@
3.5 QASs From Quantum Advantage Assumption
Definition 3.10 (Quantum Advantage Assumption [AA11, BMS16, KT24b]).
We say that quantum advantage assumption holds if the following is satisfied.
  1. 1.
    There exists a family $\mathcal{C}=\{\mathcal{C}_{\lambda}\}_{\lambda\in\mathbb{N}}$ of distributions such that for each $\lambda\in\mathbb{N}$, $\mathcal{C}_{\lambda}$ is a (uniform) QPT sampleable distribution over quantum circuits $C$ that output $\lambda$-bit classical bit strings.
  2. 2.
    There exist polynomials $p$ and $\gamma$ such that:
    1. (a)
      For all sufficiently large $\lambda\in\mathbb{N}$,
      $\Pr_{\begin{subarray}{c}C\leftarrow\mathcal{C}_{\lambda}\\ x\leftarrow\{0,1\}^{\lambda}\end{subarray}}\left[\Pr[x\leftarrow C]\geq\frac{1}{p(\lambda)2^{\lambda}}\right]\geq\frac{1}{\gamma(\lambda)}.$ (106)
    2. (b)
      For any oracle $\mathcal{O}$ satisfying that for all sufficiently large $\lambda\in\mathbb{N}$,
      $\Pr_{\begin{subarray}{c}C\leftarrow\mathcal{C}_{\lambda}\\ x\leftarrow\{0,1\}^{\lambda}\end{subarray}}\left[|\mathcal{O}(C,x)-\Pr[x\leftarrow C]|\leq\frac{\Pr[x\leftarrow C]}{p(\lambda)}\right]\geq\frac{1}{\gamma(\lambda)}-\frac{1}{p(\lambda)},$ (107)
      we have that $\mathbf{P}^{\mathbf{\#P}}\subseteq\mathbf{BPP}^{\mathcal{O}}$.
Remark 3.11.
Traditionally, we required $\mathbf{P}^{\mathbf{\#P}}\subseteq\mathbf{BPP}^{\mathcal{O}}$, but instead of it, we could consider $\mathbf{P}^{\mathbf{\#P}}\subseteq\mathbf{BQP}^{\mathcal{O}}$, for example, because $\mathbf{P}^{\mathbf{\#P}}\subseteq\mathbf{BQP}^{\mathbf{NP}}$ is also unlikely.
We show that QASs can be derived from the quantum advantage assumption (Definition 3.10) plus $\mathbf{P}^{\mathbf{\#P}}\nsubseteq\mathbf{ioBPP}^{\mathbf{NP}}$.
Theorem 3.12.
If the quantum advantage assumption (Definition 3.10) holds and $\mathbf{P}^{\mathbf{\#P}}\nsubseteq\mathbf{ioBPP}^{\mathbf{NP}}$, then QASs exist.
We can directly derive the existence of QASs from the quantum advantage assumption plus $\mathbf{P}^{\mathbf{\#P}}\nsubseteq\mathbf{ioBPP}^{\mathbf{NP}}$. However, we first introduce a useful notion, which we call hardness of quantum probability estimation (QPE), and show the theorem via hardness of QPE.
Definition 3.13 (Hardness of Quantum Probability Estimation (QPE)).
We say that hardness of quantum probability estimation (QPE) holds if the following is satisfied.
  1. 1.
    There exists a family $\mathcal{D}=\{\mathcal{D}_{\lambda}\}_{\lambda\in\mathbb{N}}$ of distributions such that for each $\lambda\in\mathbb{N}$, $\mathcal{D}_{\lambda}$ is a (uniform) QPT sampleable distribution over classical bit strings.
  2. 2.
    There exists a polynomial $p$ such that for any oracle PPT algorithm $\mathcal{A}^{\mathbf{NP}}$ and for all sufficiently large $\lambda\in\mathbb{N}$,
    $\Pr_{x\leftarrow\mathcal{D}_{\lambda}}\left[|\mathcal{A}^{\mathbf{NP}}(1^{\lambda},x)-\Pr[x\leftarrow\mathcal{D}_{\lambda}]|\leq\frac{\Pr[x\leftarrow\mathcal{D}_{\lambda}]}{p(\lambda)}\right]\leq 1-\frac{1}{p(\lambda)}.$ (108)
Remark 3.14.
A similar notion was introduced in [KT24b, CGGH24, HM24]. The main difference here is that the hardness is for PPT algorithms with $\mathbf{NP}$ oracle.
Theorem 3.12 is shown by combining the following two lemmas.
Lemma 3.15 (Based on [KT24b]).
If the quantum advantage assumption holds and $\mathbf{P}^{\mathbf{\#P}}\nsubseteq\mathbf{ioBPP}^{\mathbf{NP}}$, then hardness of QPE holds.
Lemma 3.16.
If hardness of QPE holds, then QASs exist.
The proof of Lemma 3.15 is almost the same as Theorem 5.1 of [KT24b] and therefore we omit it.
In the proof of Lemma 3.16, we use the following lemma.
Lemma 3.17 (Approximate Counting [Sto83]).
For any $f\in\mathbf{\#P}$ and for any polynomial $p$, there exists an oracle PPT algorithm $\mathcal{R}^{\mathbf{NP}}$ with $\mathbf{NP}$ oracle such that for any $x$,
$\Pr\left[|\mathcal{R}^{\mathbf{NP}}(f,x)-f(x)|\leq\frac{1}{p(|x|)}\right]\geq 1-{\mathsf{negl}}(|x|).$ (109)
Now, we are ready to show Lemma 3.16.
Proof of Lemma 3.16.
Let $\mathcal{D}=\{\mathcal{D}_{\lambda}\}_{\lambda\in\mathbb{N}}$ be a family of distributions that satisfies Definition 3.13. For each $\lambda\in\mathbb{N}$, let $\mathcal{D}_{\lambda}$ be a QPT samplable distribution over $\{0,1\}^{m(\lambda)}$, where $m$ is a polynomial. Then there exists a polynomial $p$ such that for any oracle PPT algorithm $\mathcal{A}^{\mathbf{NP}}$ and for all sufficiently large $\lambda\in\mathbb{N}$,
$\Pr_{x\leftarrow\mathcal{D}_{\lambda}}\left[|\mathcal{A}^{\mathbf{NP}}(1^{\lambda},x)-\Pr[x\leftarrow\mathcal{D}_{\lambda}]|\leq\frac{\Pr[x\leftarrow\mathcal{D}_{\lambda}]}{p(\lambda)}\right]\leq 1-\frac{1}{p(\lambda)}.$ (110)
Let us consider the QPT algorithm $\mathcal{Q}$ that takes $1^{\lambda}$ as input and samples $x\leftarrow\mathcal{D}_{\lambda}$. We show that $\mathcal{Q}$ is a QAS. For the sake of contradiction, assume that $\mathcal{Q}$ is not a QAS. Then, for any polynomial $q$, there exists a PPT algorithm $\mathcal{S}_{q}$ such that
$\mathsf{SD}(\mathcal{Q}(1^{\lambda}),\mathcal{S}_{q}(1^{\lambda}))\leq\frac{1}{q(\lambda)}$ (111)
holds for infinitely many $\lambda\in\mathbb{N}$. By Lemma 3.17, for any polynomial $p$, there exists an oracle PPT algorithm $\mathcal{A}^{\mathbf{NP}}_{p}$ such that for all $x\in\{0,1\}^{m(\lambda)}$,
$\Pr\left[|\mathcal{A}^{\mathbf{NP}}_{p}(x)-\Pr[x\leftarrow\mathcal{S}_{q}(1^{\lambda})]|\leq\frac{\Pr[x\leftarrow\mathcal{S}_{q}(1^{\lambda})]}{3p(\lambda)}\right]\geq 1-{\mathsf{negl}}(\lambda).$ (112)
In the following, we set $q(\lambda):=12p(\lambda)^{2}$. For each $\lambda\in\mathbb{N}$, we define a set $\mathsf{Good}_{\lambda}$ as follows:
$\mathsf{Good}_{\lambda}:=\left\{x\in\{0,1\}^{m(\lambda)}:|\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]-\Pr[x\leftarrow\mathcal{S}_{q}(1^{\lambda})]|\leq\frac{\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]}{3p(\lambda)}\right\}.$ (113)
Then,
$\Pr_{x\leftarrow\mathcal{Q}(1^{\lambda})}[x\notin\mathsf{Good}_{\lambda}]$ $=\Pr_{x\leftarrow\mathcal{Q}(1^{\lambda})}\left[|\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]-\Pr[x\leftarrow\mathcal{S}_{q}(1^{\lambda})]|\geq\frac{\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]}{3p(\lambda)}\right]$ (114)
$=\Pr_{x\leftarrow\mathcal{Q}(1^{\lambda})}\left[\left|1-\frac{\Pr[x\leftarrow\mathcal{S}_{q}(1^{\lambda})]}{\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]}\right|\geq\frac{1}{3p(\lambda)}\right]$ (115)
[…]
$=3p(\lambda)\sum_{x}|\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]-\Pr[x\leftarrow\mathcal{S}_{q}(1^{\lambda})]|$ (117)
$=6p(\lambda)\mathsf{SD}(\mathcal{Q}(1^{\lambda}),\mathcal{S}_{q}(1^{\lambda}))$ (118)
$\leq\frac{6p(\lambda)}{q(\lambda)}=\frac{1}{2p(\lambda)}$ (119)
holds for infinitely many $\lambda\in\mathbb{N}$. Moreover for all sufficiently large $\lambda$ and for all $x\in\mathsf{Good}_{\lambda}$,
$1-{\mathsf{negl}}(\lambda)$ (120)
$\leq\Pr\left[|\mathcal{A}^{\mathbf{NP}}_{p}(x)-\Pr[x\leftarrow\mathcal{S}_{q}(1^{\lambda})]|\leq\frac{\Pr[x\leftarrow\mathcal{S}_{q}(1^{\lambda})]}{3p(\lambda)}\right]$ (121)
$\leq\Pr\left[|\mathcal{A}^{\mathbf{NP}}_{p}(x)-\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]|\leq\frac{\Pr[x\leftarrow\mathcal{S}_{q}(1^{\lambda})]}{3p(\lambda)}+|\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]-\Pr[x\leftarrow\mathcal{S}_{q}(1^{\lambda})]|\right]$ (122)
$\leq\Pr\left[|\mathcal{A}^{\mathbf{NP}}_{p}(x)-\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]|\leq\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]\left\{\frac{1}{3p(\lambda)}\left(1+\frac{1}{3p(\lambda)}\right)+\frac{1}{3p(\lambda)}\right\}\right]$ (123)
$\leq\Pr\left[|\mathcal{A}^{\mathbf{NP}}_{p}(x)-\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]|\leq\frac{\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]}{p(\lambda)}\right].$ (124)
Therefore,
$\Pr_{x\leftarrow\mathcal{Q}(1^{\lambda})}\left[|\mathcal{A}^{\mathbf{NP}}_{p}(x)-\Pr[x\leftarrow\mathcal{D}_{\lambda}]|\leq\frac{\Pr[x\leftarrow\mathcal{D}_{\lambda}]}{p(\lambda)}\right]$ (125)
$\geq\Pr_{x\leftarrow\mathcal{Q}(1^{\lambda})}[x\in\mathsf{Good}_{\lambda}]\Pr_{x\leftarrow\mathcal{Q}(1^{\lambda})}\left[|\mathcal{A}^{\mathbf{NP}}_{p}(x)-\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]|\leq\frac{\Pr[x\leftarrow\mathcal{Q}(1^{\lambda})]}{p(\lambda)}~{}\middle|~{}x\in\mathsf{Good}_{\lambda}\right]$ (126)
$\geq\Pr_{x\leftarrow\mathcal{Q}(1^{\lambda})}[x\in\mathsf{Good}_{\lambda}](1-{\mathsf{negl}}(\lambda))$ (127)
$\geq\left(1-\frac{1}{2p(\lambda)}\right)(1-{\mathsf{negl}}(\lambda))$ (128)
$\geq 1-\frac{1}{p(\lambda)}$ (129)
holds for infinitely many $\lambda\in\mathbb{N}$. This contradicts Equation 110.
@@ 5 Equivalence of IV-PoQ and Classically-Secure OWPuzzs
For this specific $\{f_{\lambda}\}_{\lambda\in\mathbb{N}}$, there exists a PPT algorithm $\mathcal{R}_{p}$ such that Equations 151 and 152 hold for infinitely many $\lambda\in\mathbb{N}$. We write $\Lambda\subseteq\mathbb{N}$ to mean the set of such $\lambda$. Rewriting Equation 152, we have
@@
As we have stated above, for this specific $\{f_{p,\lambda}\}_{\lambda\in\mathbb{N}}$, there exists a PPT algorithm $\mathcal{R}_{p}^{f_{p,\lambda}}$ such that Equations 249 and 250 hold for infinitely many $\lambda\in\mathbb{N}$. From $\mathcal{S}_{p}$ and $\mathcal{R}_{p}^{f_{p,\lambda}}$, we construct a PPT algorithm $\mathcal{A}$ such that
@@
holds for infinitely many $\lambda\in\mathbb{N}$. Here, we have used Equation 249 to obtain Equations 258 and 264, and used Equation 250 to obtain Equation 262. Therefore, $\mathcal{A}$ breaks the soundness of classically-secure OWPuzz $(\mathsf{Samp},\mathsf{Ver})$.
@@ 6.2 Zero-Knowledge IV-PoQ
We also do not have solution for the latter approach either, but we have the following observation. We observe that we can construct statistical zero-knowledge IV-PoQ (or even efficiently-verifiable PoQ) if we additionally assume the existence of an $\mathbf{NP}$ search problem that is easy for QPT algorithms but hard for PPT algorithms (or equivalently publicly-verifiable one-round PoQ). To see this, we can consider a protocol where the honest quantum prover solves the $\mathbf{NP}$ search problem and then proves the knowledge of the solution by using statistical zero-knowledge arguments of knowledge for $\mathbf{NP}$, which exists if OWFs exist [HNO+09]. Examples of classically-hard and quantumly-easy $\mathbf{NP}$ search problems are the factoring and discrete-logarithm problems (assuming classical hardness of them) [Sho94]. Another example based on a random oracle was recently found in [YZ24]. Thus, based on the random oracle heuristic [BR93], we have a candidate construction of statistical zero-knowledge PoQ from hash functions. [^37] This is not a construction in the quantum random oracle model since we use the hash function in a non-black-box manner. Instead, we rely on the assumption that the problem considered in [YZ24] is classically hard when the random oracle is instantiated with a concrete hash function. Though this is far from a construction solely based on OWFs, this can be seen as an evidence that “structured” assumptions are not necessary for statistical zero-knowledge PoQ, let alone for statistical zero-knowledge IV-PoQ.383838“Structure” is a commonly used informal term that refers to problems behind constructions of existing public key encryption such as the hardness of factoring, discrete-logarithm, learning with errors, etc. On the other hand, hash functions are often regarded as “unstructured”. See [Bar17] for more context.
@@ Appendix A On Uniformity of Adversaries
The time to compute $f^{\prime}(z)$ is $O(|z|^{2})$ because finding the maximum $i$ spends $O(|z|^{2})$ time404040 A single-tape Turing machine spends $O(|z|^{2})$ time to find the maximum $i$. and computing $f(z_{1,...,i})$ spends $O(|z_{1,...,i}|^{c})\leq O(|z|)$ time.